### v1

# Universally Robust Control of Open Quantum Systems

Abstract
Mitigating noise-induced decoherence is the central challenge in controlling open quantum systems. While existing robust protocols often require precise noise models, we introduce a universal framework for noise-agnostic quantum control that achieves high-fidelity operations without prior environmental noise characterization. This framework capitalizes on the dynamical modification of the system-environment coupling through control drives, an effect rigorously encoded in the dynamical equation. Since the derived noise sensitivity metric remains independent of the coupling details between the system and the environment, our protocol demonstrates provable robustness against arbitrary Markovian noises. Numerical validation through quantum state transfer and gate operations reveals near-unity fidelity ($99\%$) across diverse noise regimes, achieving orders-of-magnitude error suppression compared to target-only approaches. This framework bridges critical gaps between theoretical control design and experimental constraints, establishing a hardware-agnostic pathway toward fault-tolerant quantum technologies across platforms such as superconducting circuits, trapped ions, and solid-state qubits.
I Introduction
Quantum computing and quantum information processing promise revolutionary advances in computing power Bharti et al. (2022), cryptography Portmann and Renner (2022), and sensing Degen et al. (2017). However, the fundamental challenge of maintaining quantum coherence in the face of environmental noise remains the primary obstacle to practical implementation Suter and Álvarez (2016). At the heart of these transformative technologies lies the fundamental challenge of precise manipulation of quantum states, where the fragile coherence of quantum systems must be meticulously preserved and controlled Saurav and Lidar (2025). Although effective for isolated systems, traditional quantum control strategies often become inadequate for open systems due to unmodeled system-environment interactions. Consequently, developing robust optimal control theories for open quantum systems has emerged as a critical frontier, essential for bridging the gap between idealized theoretical models and practical implementations under noisy, real-world conditions Kallush et al. (2022); Weidner et al. (2025); Gautier et al. (2025).
Optimal control in open quantum systems must reconcile two competing objectives: driving the system toward a target state while simultaneously suppressing unwanted environmental couplings Weidner et al. (2025). To date, several control protocols—including dynamical decoupling Viola et al. (1999); Ryan et al. (2010), decoherence-free subspaces Lidar et al. (1998); Bacon et al. (2000), filter-function optimization Paz-Silva and Viola (2014); Cerfontaine et al. (2021), and quantum error correction Knill et al. (2000); Beale et al. (2018)—have demonstrated significant success in this endeavor. However, their effectiveness critically depends on prior knowledge of the mathematical structure of environmental noise Gordon et al. (2008); Schulte-Herbrüggen et al. (2011); Jing et al. (2013); Maloney et al. (2022); Liu et al. (2024); Chen et al. (2025), a requirement that is rarely satisfied in experimental settings due to the inherent difficulties of comprehensive noise characterization. This gap necessitates the development of universally applicable control strategies that remain effective without detailed knowledge of system-environment coupling.
Here, we introduce a universally robust quantum control framework that substantially enhances operational fidelity in open quantum systems without requiring prior knowledge of noise characteristics. Our approach is built upon a key physical insight: external control drives not only manipulate the target system but also dynamically reconfigure the system-environment interaction itself Dann et al. (2018); Kallush et al. (2022); Di Meglio et al. (2024); Dann (2025). This dual effect is rigorously captured through a time-dependent master equation formalism, where the control fields instantaneously modulate the dissipative rates. Crucially, this description remains valid beyond adiabatic approximations and maintains thermodynamic consistency Kallush et al. (2022), establishing a physically grounded foundation for quantum control design. Through an analytical treatment of the master equation’s evolution operator, we derive an effective noise sensitivity metric that quantifies environmental susceptibility. This metric is then incorporated into a minimized objective functional defined exclusively by task-specific system observables, creating an inherently noise-agnostic optimization landscape. The resulting formulation provides provable robustness against arbitrary Markovian noises without requiring microscopic noise models—a mathematical manifestation of our protocol’s universality. We demonstrate the efficacy of our protocol through simulations of quantum state transfer and gate operations, optimized using the Chopped Random Basis (CRAB) algorithm Doria et al. (2011); Caneva et al. (2011); Müller et al. (2022). Both implementations achieve significant performance gains, obatining greater than 99% fidelity across a broad range of system-bath coupling strengths. Our framework’s platform-agnostic nature makes it broadly applicable to leading quantum hardware, including superconducting circuits Abdelhafez et al. (2020); Werninghaus et al. (2021), trapped ions Nebendahl et al. (2009); Shapira et al. (2018), and nitrogen-vacancy centers Poggiali et al. (2018); Vetter et al. (2024). This unified approach establishes a concrete pathway toward fault-tolerant quantum technologies, effectively bridging the divide between theoretical robustness and experimental feasibility.
II Results
II.1 Model
The complete quantum description of a control task is encoded in the composite Hamiltonian
$\hat{H}(t)=\hat{H}_{\rm S}(t)+\hat{H}_{\rm B}+\hat{H}_{\rm I}\,$, (1)
where $\hat{H}_{\rm S}(t)=\hat{H}_{0}+\sum_{i}u_{i}(t)\hat{H}_{i}$ represents the noise-free system Hamiltonian, with $\hat{H}_{0}$ denoting the bare drift Hamiltonian, $\hat{H}_{i}$ the $i$-th control operator, and $u_{i}(t)$ the corresponding time-dependent control field. The bath Hamiltonian $\hat{H}_{\rm B}$ governs the environment’s free dynamics, which couples to the system via interaction Hamiltonian $\hat{H}_{\rm I}=g\sum_{\alpha}\hat{A}_{\alpha}\otimes\hat{B}_{\alpha}$ (the most general form), where $\hat{A}_{\alpha}$ and $\hat{B}_{\alpha}$ are Hermitian operators of the system and the bath, respectively, while $g$ characterizes their coupling strength. Under the Born-Markov approximation, tracing over the bath degrees of freedom yields the Gorini-Kossakowski-Lindblad-Sudarshan (GKLS) master equation Lindblad (1975); Gorini et al. (1976); Lindblad (1976) (in units where $\hbar=1$; see the Appendix for a detailed derivation)
$\frac{{\rm d}}{{\rm d}t}\hat{\rho}_{\rm S}(t)=-{\rm i}[\hat{H}_{\rm S}(t),\hat{\rho}_{\rm S}(t)]+\mathcal{D}[\hat{\rho}_{\rm S}(t)]\,$, (2)
where $\hat{\rho}_{\rm S}(t)$ is the density operator of the system. The dissipative superoperator $\mathcal{D}$ takes the form
$\mathcal{D}[\hat{\rho}_{\rm S}]=\lambda\sum_{j}\kappa_{j}(t)\left[\hat{F}_{j}(t)\rho_{\rm S}(t)\hat{F}_{j}^{\dagger}(t)-\frac{1}{2}\left\{\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t),\rho_{\rm S}(t)\right\}\right]\,$, (3)
with $\lambda=g^{2}$. The Lindblad jump operator $\hat{F}_{j}(t)$ constitute the eigenoperators of the free dynamical map $\mathcal{U}_{\rm S}(t)$, satisfying
$\mathcal{U}_{\rm S}(t)\hat{F}_{j}(t)=\hat{U}_{\rm S}^{\dagger}(t)\hat{F}_{j}(t)\hat{U}_{\rm S}(t)={\rm e}^{{\rm i}\phi_{j}(t)}\hat{F}_{j}(t)\,$, (4)
where $\hat{U}_{\rm S}(t)$ is the system propagator evolving under ${\rm i}\partial_{t}\hat{U}_{\rm S}(t)=\hat{H}_{\rm S}(t)\hat{U}_{\rm S}(t)$ with initial condition $\hat{U}_{\rm S}(0)=\hat{\mathbb{I}}_{\rm S}$ (the identity operator for the system Hilbert space). The time-dependent decoherence rates $\kappa_{j}(t)$ read
$\kappa_{j}(t)=\sum_{\alpha}\left[\eta_{j}^{\alpha}(t)\right]^{2}\gamma_{\alpha\alpha}\left[\omega_{j}(t)\right]\,$, (5)
where $\eta_{j}^{\alpha}(t)=\left|\operatorname{\textnormal{Tr}}\left[{\hat{F}_{j}^{\dagger}(t)\hat{A}_{\alpha}}\right]\right|$ and $\gamma_{\alpha\alpha}\left[\omega_{j}(t)\right]$ is a function of the effective instantaneous Bohr frequencies $\omega_{j}(t)\left[={\rm d}\phi_{j}(t)/{\rm d}t\right]$.
Critically, the time-dependent master equation [Eq. (2)] explicitly accounts for how control drives in $\hat{H}_{\rm S}(t)$ modulate the dissipative dynamics. This results in control-dependent jump operators $\hat{F}_{j}(t)$ and time-varying decoherence rates $\kappa_{j}(t)$ Kallush et al. (2022). Consequently, the controller influences the system through two distinct channels: (i) Direct unitary steering via the coherent evolution ${\rm i}[\hat{H}_{\rm S}(t),\hat{\rho}_{\rm S}(t)]$ and (ii) Indirect dissipative control through the modulated dissipator $\mathcal{D}[\hat{\rho}_{\rm S}(t)]$. Remarkably, this framework enables the discovery of optimal controls $\hat{H}_{\rm S}(t)$ that simultaneously achieve target operations while actively suppressing environment-induced noise. As demonstrated in the subsequent section, such robust control protocols can be identified—even without detailed knowledge of system-bath couplings—through appropriate construction of the objective functional.
II.2 Universally Robust Noise Mitigation
We now formulate a universal approach to suppress arbitrary Markovian noise by constructing a control-dependent cost function based on the time-dependent master equation [Eq. (2)]. This cost function depends solely on the designed control Hamiltonian $\hat{H}_{\rm S}(t)$ while remaining independent of specific noise channels.
For the sake of convenience, we first vectorize the master equation Eq. (2) by rewriting it in the Hilbert-Schmidt space Caves (1999); Andersson et al. (2007). This can be down by reshaping the density matrix $\hat{\rho}_{\rm S}(t)$ as a column vector, denoted as $|\hat{\rho}_{\rm S}(t)\rangle\!\rangle$. It is straightforward to show that $|\hat{\rho}_{\rm S}(t)\rangle\!\rangle$ satisfies the following Schrödinger-type equation:
$\frac{\partial|\hat{\rho}_{\rm S}(t)\rangle\!\rangle}{\partial t}=\vec{\mathcal{L}}(t)|\hat{\rho}_{\rm S}(t)\rangle\!\rangle\,$, (6)
where the Lindbladian superoperator is vectorizd as
$\vec{\mathcal{L}}(t)=$ $-{\rm i}\left[\hat{H}_{\rm S}(t)\otimes\hat{\mathbb{I}}_{\rm S}-\hat{\mathbb{I}}_{\rm S}\otimes\hat{H}_{\rm S}(t)^{\mathrm{T}}\right]$ (7)
$+\lambda\sum_{j}\kappa_{j}(t)\left(\hat{F}_{j}(t)\otimes\hat{F}_{j}^{\ast}(t)-\frac{1}{2}\left[\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t)\otimes\hat{\mathbb{I}}_{\rm S}+\hat{\mathbb{I}}_{\rm S}\otimes\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}^{\ast}(t)\right]\right)\,.$
The advantage of this vectorization technology is that it circumvents the complexity caused by the abstract superoperator and thus facilitates the perturbation analysis we applied subsequently.
By defining the shorthand $\mathbf{H}(t)=-{\rm i}[\hat{H}_{\rm S}(t)\otimes\hat{\mathbb{I}}_{\rm S}-\hat{\mathbb{I}}_{\rm S}\otimes\hat{H}^{\rm T}_{\rm S}(t)]$ and $\mathbf{F}^{j}(t)=\hat{F}_{j}(t)\otimes\hat{F}_{j}^{\ast}(t)-\frac{1}{2}[\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t)\otimes\hat{\mathbb{I}}_{\rm S}+\hat{\mathbb{I}}_{\rm S}\otimes\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}^{\ast}(t)]$, the formal solution of Eq. (6) can be expressed as $|\hat{\rho}_{\rm S}(t)\rangle\!\rangle=\mathbf{V}(t)|\hat{\rho}(0)\rangle\!\rangle$. Here,
$\mathbf{V}(t)=\mathcal{T}\exp\left(\int_{0}^{t}{\rm d}t_{1}\left[\mathbf{H}(t_{1})+\lambda\sum_{j}\kappa_{j}(t_{1})\mathbf{F}^{j}(t_{1})\right]\right)$ (8)
is the time-evolution operator with $\mathcal{T}$ the chronological time-ordering operator. To isolate the noise effects, we transform into the interaction picture with respect to the noise-free Hamiltonian $\hat{H}_{\rm S}(t)$ and separate the evolution operator into a unitary part and an error part,
$\mathbf{V}(t)=\mathbf{U}(t)\mathbf{U}_{\mathrm{err}}(t)\,$, (9)
where $\mathbf{U}(t)=\mathcal{T}\exp\left[\int_{0}^{t}{\rm d}t_{1}\mathbf{H}(t_{1})\right]$ accounts for unitary evolution and any noise-induced non-unitary effects are attributed to the operator $\mathbf{U}_{\mathrm{err}}(t)=\mathcal{T}\exp\left[\lambda\int_{0}^{t}{\rm d}t_{1}\sum_{j}\kappa_{j}(t_{1})\widetilde{\mathbf{F}}^{j}(t_{1})\right]$ with $\widetilde{\mathbf{F}}^{j}(t)=\mathbf{U}^{\dagger}(t)\mathbf{F}^{j}(t)\mathbf{U}(t)$.
Under the assumption of weak system-bath coupling, we can expand the evolution operator $\mathbf{U}_{\mathrm{err}}(t)$ in terms of $\lambda$, leading to
$\mathbf{U}_{\mathrm{err}}(t)=\hat{\mathbb{I}}_{\rm S}\otimes\hat{\mathbb{I}}_{\rm S}$ $+\lambda\int_{0}^{t}{\rm d}t_{1}\sum_{j}\kappa_{j}(t_{1})\widetilde{\mathbf{F}}^{j}(t_{1})$ (10)
$+\lambda^{2}\int_{0}^{t}{\rm d}t_{1}\int_{0}^{t_{1}}{\rm d}t_{2}\sum_{j_{1},\,j_{2}}\kappa_{j_{1}}(t_{1})\kappa_{j_{2}}(t_{2})\widetilde{\mathbf{F}}^{j_{1}}(t_{1})\widetilde{\mathbf{F}}^{j_{2}}(t_{2})+\cdots\,.$
Robust quantum control requires insensitivity to first-order effects in the system-bath coupling strength $\lambda$. We therefore focus on the leading-order contribution, namely,
$\left.\frac{{\rm d}\mathbf{U}_{\mathrm{err}}(t)}{{\rm d}\lambda}\right|_{\lambda=0}=\int_{0}^{t}{\rm d}t_{1}\sum_{j}\kappa_{j}(t_{1})\widetilde{\mathbf{F}}^{j}(t_{1})=\int_{0}^{t}{\rm d}t_{1}\sum_{\alpha,\,j}\left[\eta_{j}^{\alpha}(t_{1})\right]^{2}\gamma_{\alpha\alpha}\left[\omega_{j}(t_{1})\right]\widetilde{\mathbf{F}}^{j}(t_{1})\,.$ (11)
The norm of Eq. (11) then quantifies the noise sensitivity, the minimization of which amounts to increasing the robustness of a control protocol. Aiming at acquiring a universal robustness, we first apply the Cauchy-Schwarz inequality to $\eta_{j}^{\alpha}(t_{1})$, yielding
$\left[\eta_{j}^{\alpha}(t_{1})\right]^{2}=\left|\operatorname{\textnormal{Tr}}\left[{\hat{F}_{j}^{\dagger}(t_{1})\hat{A}_{\alpha}}\right]\right|^{2}\leq\operatorname{\textnormal{Tr}}\left[{\hat{F}_{j}^{\dagger}(t_{1})\hat{F}_{j}(t_{1})}\right]\cdot\operatorname{\textnormal{Tr}}\left[{\hat{A}_{\alpha}^{\dagger}\hat{A}_{\alpha}}\right]=\operatorname{\textnormal{Tr}}\left[{\hat{A}_{\alpha}^{\dagger}\hat{A}_{\alpha}}\right]\,.$ (12)
Substituting Eq. (12) into Eq. (11), we have
[…]
where $\mathbf{F}_{\alpha}=\int_{0}^{t}{\rm d}t_{1}\sum_{j}\gamma_{\alpha\alpha}\left[\omega_{j}(t_{1})\right]\widetilde{\mathbf{F}}^{j}(t_{1})$ and $\left\lVert\cdot\right\rVert$ denotes the Frobenius norm. In deriving the last inequality in Eq. (13), we have made use of the Cauchy-Schwarz inequality twice. Equation (13) suggests an effective noise sensitivity $D_{\rm eff}$, defined as
$D_{\rm eff}=\sqrt{\sum_{\alpha}\left\lVert\mathbf{F}_{\alpha}\right\rVert^{2}}\,.$ (14)
It can be seen clearly that the goal of increasing the robustness of a control problem comes down to finding a $\hat{H}_{\rm S}(t)$ to minimize $D_{\rm eff}$. We emphasize that this noise mitigation scenario is quite universal in the sense that the effective noise sensitivity $D_{\rm eff}$ is independent of the specific system-bath coupling operators $\hat{A}_{\alpha}$ and is therefore effective in all Markovian environments. Moreover, while we here focus on the first-order contribution of the noise perturbation, higher-order expansion terms from Eq. (10) could in principle be incorporated into $D_{\rm eff}$ to further enhance robustness.
II.3 Optimal Control Framework
We now demonstrate how our universal robustness measure $D_{\rm eff}$ integrates with quantum optimal control. The primary objective is to implement a target unitary operation $\hat{U}_{\rm tar}$ by optimizing the control fields $u_{i}(t)$ in the system Hamiltonian $\hat{H}_{\rm S}(t)$ [see details below Eq. (1)]. This is achieved by maximizing fidelity between $\hat{U}_{\rm tar}$ and the realized evolution $\hat{U}_{\rm S}(\tau)$, where $\tau$ is total evolution time. We distinguish two control paradigms: (i) For the task of pure state transfer from an initial state $\hat{\rho}_{\rm i}$ to a target state $\hat{\rho}_{\rm tar}$, such fidelity reduces to
$\mathcal{F}_{\text{state}}=\operatorname{\textnormal{Tr}}\left[{\hat{\rho}_{\rm f}\hat{\rho}_{\rm tar}}\right]\,.$ (15)
where $\hat{\rho}_{\rm f}=\hat{U}_{\rm S}(\tau)\hat{\rho}_{\rm i}\hat{U}^{\dagger}_{\rm S}(\tau)$ and $\hat{\rho}_{\rm tar}=\hat{U}_{\rm tar}\hat{\rho}_{\rm i}\hat{U}^{{\dagger}}_{\rm tar}$. (ii) For quantum gate operations, on the other hand, the fidelity is formulated in terms of a complete set of pure initial states $\{\hat{\rho}_{\rm i}^{n}\}$ as
$\mathcal{F}_{\text{gate}}=\frac{1}{N^{2}-1}\sum_{n=1}^{N^{2}-1}\operatorname{\textnormal{Tr}}\left[{\hat{\rho}_{\rm f}^{n}\hat{\rho}_{\rm tar}^{n}}\right]\,$, (16)
where $\hat{\rho}_{\rm f}^{n}=\hat{U}_{\rm S}(\tau)\hat{\rho}_{\rm i}^{n}\hat{U}^{\dagger}_{\rm S}(\tau)$, $\hat{\rho}_{\rm tar}^{n}=\hat{U}_{\rm tar}\hat{\rho}_{\rm i}^{n}\hat{U}^{{\dagger}}_{\rm tar}$, and $N$ is the dimension of system’s Hilbert space. Note that we have excluded the identity since it is preserved in the CPTP map and this definition guarantees $\mathcal{F}_{\text{gate}}\leq 1$. In the following, we optimize the infidelity $\mathcal{J}_{0}=1-\mathcal{F}_{\rm state/gate}$ for numerical efficiency, converting fidelity maximization to minimization of a positive definite functional.
Robust optimal control additionally requires that the control process owns immunity to environment noise. We then face a multi-objective optimization task in which the total objective functional to be minimized can be organized as
$\mathcal{J}=\mathcal{J}_{0}+cD_{\rm eff}\,$, (17)
where $c$ is a weighting coefficient that balances operational accuracy ($\mathcal{J}_{0}$) against noise robustness ($D_{\rm eff}$). This leads to two distinct control strategies:
  • Target-only control ($c=0$): Optimizes for fidelity alone.
  • Universally robust control ($c>0$): Jointly optimizes for fidelity and noise suppression.
We implement both control strategies using the CRAB algorithm Doria et al. (2011); Caneva et al. (2011); Müller et al. (2022), which offers efficient optimization with experimentally advantageous features like inherent pulse smoothness and a reduced parameter space. The specific parameterization is $u_{i}(t)=\exp(-[(t-\tau/2)/(2\sigma)]^{2})\sum_{k=1}^{M}c_{k}\sin(\nu_{k}t)$, where $c_{k}$} are optimization coefficients, $\sigma$ controls the Gaussian envelope width, and $\nu_{k}=k\pi/\tau$ ($k=1,\dots,M$) are fixed frequencies (in the following, we choose $\sigma=\tau/4$ and $M=10$). We optimize $c_{k}$} using a quasi-Newton algorithm to minimize $\mathcal{J}$.
Figure 1: Universally robust quantum state transfer in a two-level system. Quantum state transfer from initial state $\hat{\rho}_{-x}$ [Bloch vector (1,0,0)(-1,0,0)] to target state $\hat{\rho}_{+x}$ [Bloch vector $(+1,0,0)$] under environmental noise. Gray dash-dotted lines: target-only control ($c=0$). Orange solid lines: universally robust control ($c=10^{-2}$). (a) State fidelity $\mathcal{F}_{\rm state}$ versus coupling strength $\lambda$ for specific noise ($\hat{A}=\hat{\sigma}_{z}$). Inset: Log-log infidelity ($1-\mathcal{F}_{\rm state}$). (b) $\mathcal{F}_{\rm state}$ versus $\lambda$ for generic noise ($\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$; $\mathbf{n}$: random unit vector). Data averaged over 20 realizations. Inset: Corresponding infidelity. (c) Optimized control fields $u_{x}(t)$ and $u_{y}(t)$ for $\hat{A}=\hat{\sigma}_{z}$ and $\lambda=0.1$. (d) State evolution trajectories on Bloch sphere for $\hat{A}=\hat{\sigma}_{z}$ and $\lambda=0.1$. Initial state: black square. Final states: blue dot (target-only) and green diamond (universally robust control). Parameters: Energy splitting $\Delta=3\times 10^{-3}$ a.u., inverse temperature $\beta=1/\Delta$.
II.4 Performance Evaluation of Universally Robust Control
To validate our framework, we implement the universally robust control protocol for two fundamental quantum tasks: (i) State transfer and (ii) Quantum gate operations under environmental noise. In the simulations of these control tasks, we model the environment as a thermal bath of harmonic oscillators with $\hat{H}_{\rm B}=\sum_{k}\left(\omega_{k}\hat{a}_{k}^{\dagger}\hat{a}_{k}+\frac{1}{2}\right)$, where $\hat{a}_{k}$ are bosonic annihilation operators for mode $k$ Breuer and Petruccione (2002). The system-environment coupling has the generic form $\hat{H}_{\rm I}=g\hat{A}\otimes\hat{B}$, where $\hat{B}=\sum_{k}\left(g_{k}/g\right)(\hat{a}_{k}+\hat{a}_{k}^{\dagger})$ and a super-Ohmic spectral density $J(\omega)=\omega_{\rm c}^{-2}\omega^{3}{\rm e}^{-\omega/\omega_{\rm c}}$ is assumed Weiss (2021). Here, $\omega_{\rm c}$ is the cutoff frequency, chosen as 10 times of the typical Bohr frequency of the bare system. Such a parameters choice corresponds to a large class of quantum systems Alkauskas et al. (2014); Jahnke et al. (2015); Norambuena et al. (2016). Besides, all calculations employ atomic units (a.u.) in the following.
II.4.1 State Transfer in a Two-Level System
We first demonstrate our quantum control protocol for state transfer implementation in a driven two-level system governed by the Hamiltonian Liu et al. (2024):
$\hat{H}_{\rm S}(t)=\frac{\Delta}{2}\hat{\sigma}_{z}+\frac{u_{x}(t)}{2}\hat{\sigma}_{x}+\frac{u_{y}(t)}{2}\hat{\sigma}_{y}\,$, (18)
where $\hat{\sigma}_{i}$ ($i=x,y,z$) denote Pauli matrices, $\Delta$ represents the energy splitting, and $u_{x,\,y}(t)$ correspond to independent control fields. This system exhibits complete controllability as the drift Hamiltonian ($\hat{\sigma}_{z}$) and the control Hamiltonians ($\hat{\sigma}_{x,\,y}$) generate the full $\mathfrak{su}(2)$ Lie algebra Huang et al. (1983); Kallush and Kosloff (2011).
We implement quantum state transfer between orthogonal maximally coherent states: initial state $\hat{\rho}_{-x}=\frac{1}{2}(\hat{\mathbb{I}}-\hat{\sigma}_{x})$ [Bloch vector (1,0,0)(-1,0,0)] and target state $\hat{\rho}_{+x}=\frac{1}{2}(\hat{\mathbb{I}}+\hat{\sigma}_{x})$ [Bloch vector $(+1,0,0)$]. To evaluate protocol robustness, we consider two distinct noise channels: (i) specific coupling $\hat{A}=\hat{\sigma}_{z}$ and (ii) generic stochastic coupling $\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$ with $\mathbf{n}$ being a randomly oriented unit vector. Control fields were optimized by minimizing the objective functional $\mathcal{J}$ [Eq. (17)] via the CRAB algorithm, comparing target-only ($c=0$) versus robust ($c>0$) control strategies. Crucially, the identical control fields optimized for $\hat{A}=\hat{\sigma}_{z}$ were applied to the generic coupling case $\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$ without re-optimization to demonstrate universality.
Fig. 1 highlights the striking advantages of our universally robust protocol. For a specific noise channel $\hat{A}=\hat{\sigma}_{z}$ [Fig. 1 (a)], both protocols achieve near-unity fidelity ($\mathcal{F}_{\text{state}}\approx 1$) at zero system-bath coupling ($\lambda=0$). However, under increasing $\lambda$, the target-only optimization exhibits rapid fidelity degradation $\mathcal{F}_{\text{state}}<0.7$ at $\lambda=0.1$], while the robust protocol maintains $\mathcal{F}_{\text{state}}>0.99$ throughout $\lambda\in[0,0.1]$. More importantly, when faced with generic, uncharacterized noise $\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$ [Fig. 1 (b)], the robust control preserves high fidelity ($\mathcal{F}_{\text{state}}>0.99$ at $\lambda=0.1$), whereas the target-only approach fails catastrophically ($\mathcal{F}_{\text{state}}<0.7$). This clearly confirms the universality of our robust control protocol. Beyond its superior fidelity, the robust protocol is also significantly more efficient, requiring control field amplitudes that are approximately 50% lower than the target-only method [Fig. 1 (c)]. Furthermore, it excels at preserving quantum state purity. As illustrated by the Bloch sphere trajectories [Fig. 1 (d)], the robust evolution remains confined to the sphere’s surface, indicating a pure, unitary process, whereas the target-only path spirals inward—a clear signature of decoherence. These combined features demonstrate a capacity for universal noise resistance with minimal operational overhead.
Figure 2: Universally robust control for single-qubit Hadamard gate. Gray dash-dotted lines: target-only control ($c=0$). Red solid lines: universally robust control ($c=10^{-2}$). (a) Gate fidelity $\mathcal{F}_{\rm gate}$ versus coupling strength $\lambda$ for specific noise ($\hat{A}=\hat{\sigma}_{z}$). Inset: Log-log infidelity ($1-\mathcal{F}_{\rm gate}$). (b) $\mathcal{F}_{\rm gate}$ versus $\lambda$ for generic noise ($\hat{A}=\mathbf{n}\cdot\hat{\boldsymbol{\sigma}}$; $\mathbf{n}$: random unit vector). Data averaged over 20 realizations. Inset: Corresponding infidelity. (c) Optimized control fields $u_{x}(t)$ and $u_{y}(t)$ for $\hat{A}=\hat{\sigma}_{z}$ and $\lambda=0.1$. (d) State evolution trajectories on Bloch sphere under the Hadamard gate (from the z-z to the x-x direction) for $\hat{A}=\hat{\sigma}_{z}$ and $\lambda=0.1$. Initial state: black square. Final states: blue dot (target-only) and green diamond (universally robust control). Parameters: Energy splitting $\Delta=3\times 10^{-3}$ a.u., inverse temperature $\beta=1/\Delta$.
Figure 3: Universally robust control for two-qubit CZ Gate. Gray dash-dotted lines: target-only control ($c=0$). Purple solid lines: universally robust control ($c=2\times 10^{-3}$). (a) Gate fidelity $\mathcal{F}_{\rm gate}$ versus coupling strength $\lambda$ for specific noise ($\hat{A}=\hat{\sigma}_{y}\otimes\hat{\sigma}_{0}$). Inset: Log-log infidelity ($1-\mathcal{F}_{\rm gate}$). (b) $\mathcal{F}_{\rm gate}$ versus $\lambda$ for generic noise $\hat{A}=\sum_{\mu\nu}a_{\mu\nu}\hat{\sigma}_{\mu}\otimes\hat{\sigma}_{\nu}$, where the random coefficients $a_{\mu\nu}\sim N(0,1)$ with $\sum_{\mu\nu}a^{2}_{\mu\nu}=1$]. Data averaged over 20 realizations. Inset: Corresponding infidelity. (c) Optimized control fields $u_{j}(t)$ ($j=1,2,3,4$) for $\hat{A}=\hat{\sigma}_{y}\otimes\hat{\sigma}_{0}$ and $\lambda=0.1$. Parameters: Energy splitting $\Delta=3\times 10^{-3}$ a.u., inverse temperature $\beta=1/\Delta$.
II.4.2 Quantum Gate Operations
We extend our robust control protocol to the more demanding task of quantum gate synthesis, which require noise-resilient state transformations across the entire Hilbert space—a significantly more complex challenge than single-state transfer Palao and Kosloff (2003); Aspman et al. . We demonstrate protocol effectiveness for two fundamental gates: the single-qubit Hadamard gate and two-qubit controlled-Z (CZ) gate, which form a universal gate set when combined with T gates Nielsen and Chuang (2010).
Let’s first consider the single-qubit Hadamard gate with target transformation
$\hat{U}_{\rm H}=\frac{1}{\sqrt{2}}\begin{pmatrix}1&1\\ 1&-1\end{pmatrix}\,$. (19)
Using the control Hamiltonian from Eq. (18), we observe performance trends mirroring the two-level state transfer results. The influence of the noise strength $\lambda$ on the gate fidelity with $\hat{A}=\hat{\sigma}_{z}$ and $\hat{A}={\bf n}\cdot\hat{\boldsymbol{\sigma}}$ are depicted in Fig. 2 (a) and (b), respectively. In both cases, the target-only optimization results deviate substantially from the ideal value once $\lambda\neq 0$, while those of robust control exhibit strong noise robustness (e.g., preserves $\mathcal{F}_{\text{gate}}>0.99$ at $\lambda=0.1$ [Fig. 2 (b)]). Compared to the target-only optimization, the fidelity degradation of the robust protocol is suppressed by two orders of magnitude. The comparison of the required control fields in the two protocols is shown in Fig. 2 (c), where we find again that the robust optimization requires much lower control-field amplitude than the target-only one. Besides, robust trajectories maintain confinement to the Bloch sphere surface (purity preservation), contrasting with target-only penetration [Fig. 2 (d)].
For the two-qubit CZ gate, the target transformation is given by:
$\hat{U}_{\rm CZ}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&0&0&-1\end{pmatrix}\,$. (20)
The noise-free control Hamiltonian for this gate is organized as Riaz et al. (2019)
$\hat{H}_{\rm S}(t)=\Delta\hat{\sigma}_{z}\otimes\hat{\sigma}_{z}$ $+u_{1}(t)\hat{\sigma}_{y}\otimes\hat{\sigma}_{0}+u_{2}(t)\hat{\sigma}_{0}\otimes\hat{\sigma}_{y}$ (21)
$+u_{3}(t)\hat{\sigma}_{z}\otimes\hat{\sigma}_{0}+u_{4}(t)\hat{\sigma}_{0}\otimes\hat{\sigma}_{z}\,$,
where $\hat{\sigma}_{0}$ is the identity operator for qubit. The control fields to be optimized are represented by $u_{j}(t)$ ($j=1,2,3,4$). Fig. 3 shows the variation of the gate fidelities against $\lambda$ for (a) $\hat{A}=\hat{\sigma}_{y}\otimes\hat{\sigma}_{0}$ and (b) $\hat{A}=\sum_{\mu\nu}a_{\mu\nu}\hat{\sigma}_{\mu}\otimes\hat{\sigma}_{\nu}$. Here, $\mu,\,\nu=0,x,y,z$, and the random coefficients $a_{\mu\nu}\sim N(0,1)$ with $\sum_{\mu\nu}a^{2}_{\mu\nu}=1$. In both cases, we find qualitatively the same behaviors as those of the single-qubit gate. That is, the results of the target-only optimization decreases rapidly as $\lambda$ increases, while the robust control can achieve and maintain high fidelity across a considerable range of noise amplitudes (e.g., preserves $\mathcal{F}_{\text{gate}}>0.99$ at $\lambda=0.1$ [Fig. 3 (b)]). Besides, as shown in Fig. 3 (c), the robust control requires significantly lowered control-field amplitudes. The robust protocol consistently outperforms target-only optimization across all metrics, demonstrating scalability to multi-qubit systems while maintaining noise resilience and control efficiency.
III Discussion
We have developed and validated a universally robust control framework capable of suppressing noise in open quantum systems without requiring specific knowledge of the noise channels. By engineering a control-objective functional that minimizes an intrinsic noise-susceptibility metric, our method achieves provable robustness against arbitrary Markovian noise. Numerical simulations of state transfer and quantum gate operations show that this approach yields near-unity fidelities and suppresses errors by orders of magnitude compared to target-only optimization, even when subjected to unanticipated noise sources.
This work represents a significant paradigm shift for robust quantum control. Traditionally, high-performance control has relied on methods like dynamical decoupling or filter-function engineering, which require at least partial characterization of the noise environment—a persistent experimental bottleneck. Our framework circumvents this requirement, decoupling control design from the arduous task of environmental characterization. This not only reduces experimental overhead but also provides resilience against uncharacterized or time-varying noise sources, a common challenge in real-world quantum devices.
Looking forward, several avenues for future research are apparent. The current framework is derived under the Born-Markov approximation. Extending this approach to combat non-Markovian noise, which possesses memory effects, is a crucial next step and would broaden its applicability to an even wider range of physical systems. While our method relies on a first-order expansion of the noise effects, its formulation allows for the inclusion of higher-order terms from Eq. (10). Investigating the trade-off between the enhanced robustness from higher-order corrections and the increased computational cost of optimization would be a valuable pursuit.
Furthermore, the scalability of the optimization process presents a practical consideration. Although we have demonstrated success for a two-qubit gate, the computational resources required to calculate and minimize $D_{\rm eff}$ will grow with system size. Developing more efficient numerical techniques or leveraging machine learning to navigate the vast parameter space of many-qubit systems will be essential for applying this framework to larger-scale quantum processors.
In conclusion, our universally robust control protocol provides a powerful and practical tool for mitigating decoherence in quantum technologies. Its hardware-agnostic nature and independence from noise-model specification offer a versatile solution applicable to quantum computing Berberich and Fink (2024), high-precision quantum sensing Hecht et al. (2025), and other quantum control tasks plagued by poorly characterized noise. By bridging a critical gap between theoretical control design and experimental reality, this work establishes a concrete and promising pathway toward the realization of fault-tolerant quantum information processing.
Acknowledgments
X.Q. acknowledges support from Shanghai Science and Technology project (24LZ1401600) and the Fundamental Research Funds for the Central Universities. J.F. acknowledges support from the National Key R&D Program of China under Grant No. 2022YFA1404201, the National Natural Science Foundation of China (NSFC) under Grants No. 12174233, the Research Project Supported by Shanxi Scholarship Council of China and Shanxi ‘1331KSC’.
Appendix: derivation of control master equation
We derive the time-dependent master equation Eq. (2) starting from the composite Hamiltonian $\hat{H}(t)=\hat{H}_{\rm S}(t)+\hat{H}_{\rm B}+\hat{H}_{\rm I}$ [Eq. (1)] under the Born-Markov approximation. Transforming to the interaction picture yields:
$\tilde{H}_{\rm I}(t)=g\sum_{\alpha}\tilde{A}_{\alpha}(t)\otimes\tilde{B}_{\alpha}(t)\,$, (22)
where $\tilde{A}_{\alpha}(t)=\hat{U}_{\rm S}^{\dagger}(t)\hat{A}_{\alpha}\hat{U}_{\rm S}(t)$ and $\tilde{B}_{\alpha}(t)=\hat{U}_{\rm B}^{\dagger}(t)\hat{B}_{\alpha}\hat{U}_{B}(t)$ with $\hat{U}_{\rm S}(t)=\mathcal{T}\exp[-{\rm i}\int_{0}^{t}\hat{H}_{\rm S}(s)\,ds]$ and $\hat{U}_{\rm B}(t)=\exp[-{\rm i}\hat{H}_{\rm B}t]$ the time evolution operators of the system and the bath, respectively. By expanding the Liouville-von Neumann equation to second order in the coupling strength $g$ and applying the Born-Markov approximation, one can derive the master equation Breuer and Petruccione (2002)
$\frac{{\rm d}}{{\rm d}t}\tilde{\rho}_{\rm S}(t)=-g^{2}\int_{0}^{\infty}{\rm d}s\,\text{Tr}_{\rm B}\left[\tilde{H}(t),\left[\tilde{H}(t-s),\,\tilde{\rho}_{\rm S}(t)\otimes\tilde{\rho}_{\rm B}\right]\right]\,.$ (23)
[…]
$\tilde{A}_{\alpha}(t)$ $=\hat{U}_{\rm S}^{\dagger}(t)\hat{A}_{\alpha}\hat{U}_{\rm S}(t)=\sum_{j}\hat{U}_{\rm S}^{\dagger}(t)c_{j}^{\alpha}\hat{F}_{j}(t)\hat{U}_{\rm S}(t)=\sum_{j}c_{j}^{\alpha}{\rm e}^{{\rm i}\phi_{j}(t)}\hat{F}_{j}(t)$ (24)
$=\sum_{j}\eta_{j}^{\alpha}\hat{F}_{j}(t){\rm e}^{{\rm i}\Lambda_{j}^{\alpha}(t)}=\sum_{j}\eta_{j}^{\alpha}\hat{F}_{j}^{\dagger}(t){\rm e}^{-{\rm i}\Lambda_{j}^{\alpha}(t)}\,$,
[…]
Substituting Eq. (24) into Eq. (23), we obtain
$\frac{{\rm d}}{{\rm d}t}\tilde{\rho}_{\rm S}(t)=g^{2}\sum_{\alpha\alpha^{\prime},\,jj^{\prime}}\int_{0}^{\infty}{\rm d}s\left\{\eta_{j^{\prime}}^{\alpha^{\prime}}(t)\eta_{j}^{\alpha}(t-s)\times{\rm Tr}_{\rm B}\left[\tilde{B}_{\alpha^{\prime}}(t)\tilde{B}_{\alpha}(t-s)\hat{\rho}_{\rm B}\right]{\rm e}^{{\rm i}\left[\Lambda_{j}^{\alpha}(t-s)-\Lambda_{j^{\prime}}^{\alpha^{\prime}}(t)\right]}\right$. (25)
$\left.\times\left[\hat{F}_{j}(t-s)\tilde{\rho}_{S}(t)\hat{F}_{j^{\prime}}^{\dagger}(t)-\hat{F}_{j^{\prime}}^{\dagger}(t)\hat{F}_{j}(t-s)\tilde{\rho}_{S}(t)\right]+\text{h.c.}\right\}\,$,
where h.c. denotes the hermitian conjugate. Under Markovian dynamics, the characteristic correlation time of the environment $\tau_{\rm E}$ is much shorter than the typical time of the drive, implying that the integral with respect to $s$ is dominated by the value of the integrand in the range $s\in\left[0,\tau_{\rm E}\right]$. In such a short time scale, it is reasonable to approximate $\eta_{j}^{\alpha}(t-s)\approx\eta_{j}^{\alpha}(t)$ and $\hat{F}_{j}(t-s)\approx\hat{F}_{j}(t)$, leading to
$\frac{{\rm d}}{{\rm d}t}\tilde{\rho}_{\rm S}(t)=\sum_{jj^{\prime}}\Phi_{jj^{\prime}}(t)\left[\hat{F}_{j}(t)\tilde{\rho}_{\rm S}(t)\hat{F}_{j^{\prime}}^{\dagger}(t)-\hat{F}_{j^{\prime}}^{\dagger}(t)\hat{F}_{j}(t)\tilde{\rho}_{\rm S}(t)\right]+\text{h.c.}\,$, (26)
with
$\Phi_{jj^{\prime}}(t)=g^{2}\sum_{\alpha\alpha^{\prime}}\int_{0}^{\infty}{\rm d}s\,\eta_{j^{\prime}}^{\alpha^{\prime}}(t)\eta_{j}^{\alpha}(t){\rm Tr}_{\rm B}\left[\tilde{B}_{\alpha^{\prime}}(s)\tilde{B}_{\alpha}(0)\hat{\rho}_{\rm B}\right]{\rm e}^{{\rm i}\left[\Lambda_{j}^{\alpha}(t-s)-\Lambda_{j^{\prime}}^{\alpha^{\prime}}(t)\right]}\,$, (27)
where we have used the time-translation invariance of the environmental correlations. The rapid decay of environmental correlations allows expanding the phases near $t$ Kallush et al. (2022); Wu et al. (2022), i.e.,
$\Lambda_{j}^{\alpha}(t-s)=\Lambda_{j}^{\alpha}(t)-\dot{\Lambda}_{j}^{\alpha}(t)s\approx\Lambda_{j}^{\alpha}(t)-\omega_{j}(t)s\,$, (28)
where $\omega_{j}(t)={\rm d}\phi_{j}(t)/{\rm d}t$.
The substitution of Eq. (28) into Eq. (25) gives rise to
$\frac{{\rm d}}{{\rm d}t}\tilde{\rho}_{\rm S}(t)=g^{2}\sum_{\alpha\alpha^{\prime},\,jj^{\prime}}$ $\left\{\eta_{j^{\prime}}^{\alpha^{\prime}}(t)\eta_{j}^{\alpha}(t){\rm e}^{{\rm i}\left[\Lambda_{j}^{\alpha}(t)-\Lambda_{j^{\prime}}^{\alpha^{\prime}}(t)\right]}\times\right$. (29)
$\left.\Gamma_{\alpha\alpha^{\prime}}\left[\omega_{j}(t)\right]\left[\hat{F}_{j}(t)\tilde{\rho}_{\rm S}(t)\hat{F}_{j^{\prime}}^{\dagger}(t)-\hat{F}_{j^{\prime}}^{\dagger}(t)\hat{F}_{j}(t)\tilde{\rho}_{\rm S}(t)\right]+\text{h.c.}\right\}\,$,
where $\Gamma_{\alpha\alpha^{\prime}}\left[\omega_{j}(t)\right]$ is the Fourier transform of the instantaneous bath correlation function, i.e.,
$\Gamma_{\alpha\alpha^{\prime}}\left[\omega_{j}(t)\right]=\int_{0}^{\infty}{\rm d}s\,{\rm e}^{-{\rm i}\omega_{j}(t)s}\,{\rm Tr}_{\rm B}\left[\tilde{B}_{\alpha^{\prime}}(s)\tilde{B}_{\alpha}(0)\hat{\rho}_{\rm B}\right]\,.$ (30)
Using the identity $\int_{0}^{\infty}{\rm d}s\,{\rm e}^{-{\rm i}\varepsilon s}=\pi\delta(\varepsilon)-i\mathcal{P}\frac{1}{\varepsilon}$ [here $\delta(\varepsilon)$ is the Dirac delta function and $\mathcal{P}$ is the Cauchy principle value], $\Gamma_{\alpha\alpha^{\prime}}\left[\omega_{j}(t)\right]$ can be rewritten as a sum of real and imaginary parts
$\Gamma_{\alpha\alpha^{\prime}}(\omega)=\frac{1}{2}\gamma_{\alpha\alpha^{\prime}}(\omega)+{\rm i}S_{\alpha\alpha^{\prime}}(\omega)\,$, (31)
where $\gamma_{\alpha\alpha^{\prime}}(\omega)=\int_{-\infty}^{\infty}{\rm d}s\,{\rm e}^{-{\rm i}\omega s}\,{\rm Tr}_{\rm B}\left[\tilde{B}_{\alpha^{\prime}}(s)\tilde{B}_{\alpha}(0)\hat{\rho}_{\rm B}\right]$ and $S_{\alpha\alpha^{\prime}}(\omega)=\frac{1}{2{\rm i}}\left[\Gamma_{\alpha\alpha^{\prime}}(\omega)-\Gamma_{\alpha^{\prime}\alpha}^{\ast}(\omega)\right]$.
We then perform the secular approximation which neglects fast oscillating terms with $\Lambda_{j}^{\alpha}(t)\neq\Lambda_{j^{\prime}}^{\alpha^{\prime}}(t)$ in the Master equation (29), yielding
$\frac{{\rm d}}{{\rm d}t}\tilde{\rho}_{\rm S}(t)=$ $-{\rm i}\left[\tilde{H}_{\rm LS}(t),\tilde{\rho}_{S}(t)\right]$ (32)
$+g^{2}\sum_{\alpha,\,j}\left[\eta_{j}^{\alpha}(t)\right]^{2}\gamma_{\alpha\alpha}\left[\omega_{j}(t)\right]\left[\hat{F}_{j}(t)\tilde{\rho}_{\rm S}(t)\hat{F}_{j}^{\dagger}(t)-\frac{1}{2}\left\{\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t),\tilde{\rho}_{\rm S}(t)\right\}\right]\,$,
where $\tilde{H}_{\rm LS}(t)=\sum_{\alpha,\,j}S_{\alpha\alpha}\left[\omega_{j}(t)\right]\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t)$ represents the time-dependent Lamb-type shift Hamiltonian. Transforming back to the Schrödinger picture and neglect the Lamb shift terms, we have
$\frac{{\rm d}}{{\rm d}t}\rho_{\rm S}(t)=$ $-{\rm i}\left[H_{\rm S}(t),\rho_{\rm S}(t)\right]$ (33)
$+g^{2}\sum_{\alpha,\,j}\left[\eta_{j}^{\alpha}(t)\right]^{2}\gamma_{\alpha\alpha}\left[\omega_{j}(t)\right]\left[\hat{F}_{j}(t)\rho_{\rm S}(t)\hat{F}_{j}^{\dagger}(t)-\frac{1}{2}\left\{\hat{F}_{j}^{\dagger}(t)\hat{F}_{j}(t),\rho_{\rm S}(t)\right\}\right]\,.$
While Eq. (33) is generically valid under the mentioned approximations, to facilitate the numerical implementation, it is preferable to rewrite it in a form constituted by parameters given in the Results section. This can be conveniently down by relabelling the jump operators through $\hat{F}_{j}(t)\mapsto\hat{F}_{i,+}(t)$ and $\hat{F}_{j}^{\dagger}(t)\mapsto\hat{F}_{i,-}(t)$, noting that $\hat{F}_{j}(t)$ and $\hat{F}_{j}^{\dagger}(t)$ are basically orthogonal to each other in the Hilbert-Schmidt space provided $\hat{F}_{j}(t)\neq\hat{F}_{j}^{\dagger}(t)$. Here, the subscript $i$ counts the conjugate pairs of eigenoperators, and $\pm$ specifies one of the two operators in each conjugate pair. With such reclassification of the dissipation channels, Eq. (33) can be rewritten as
$\frac{{\rm d}}{{\rm d}t}\rho_{\rm S}(t)=$ $-{\rm i}\left[H_{\rm S}(t),\rho_{\rm S}(t)\right]$ (34)
$+g^{2}\sum_{\alpha,\,i}\left[\eta_{i}^{\alpha}(t)\right]^{2}\left(\kappa_{\alpha\alpha}^{+}\left[\omega_{i}(t)\right]\left[\hat{F}_{i,+}(t)\rho_{\rm S}(t)\hat{F}_{i,+}^{\dagger}(t)-\frac{1}{2}\left\{\hat{F}_{i,+}^{\dagger}(t)\hat{F}_{i,+}(t),\rho_{\rm S}(t)\right\}\right]\right$.
$\qquad\qquad\qquad+\left.\kappa_{\alpha\alpha}^{-}\left[\omega_{i}(t)\right]\left[\hat{F}_{i,-}(t)\rho_{\rm S}(t)\hat{F}_{i,-}^{\dagger}(t)-\frac{1}{2}\left\{\hat{F}_{i,-}^{\dagger}(t)\hat{F}_{i,-}(t),\rho_{\rm S}(t)\right\}\right]\right)\,$,
where $\kappa_{\alpha\alpha}^{\pm}\left[\omega_{j}(t)\right]=\gamma_{\alpha\alpha}\left[\pm\omega_{j}(t)\right][1-\delta_{\omega_{j}(t),0}/2]$ with $\delta_{\omega_{j}(t),0}$ denoting the kronecker delta function. For the quantum tasks we considered in the Results section, the sum over the subscript $\alpha$ in Eq. (34) disappears and the kinetic coefficients reduces to $\kappa^{+}\left[\omega_{j}(t)\right]=2\pi J\left[\omega_{j}(t)\right]\left(\overline{N}\left[\omega_{j}(t)\right]+1\right)$ and $\kappa^{-}\left[\omega_{j}(t)\right]=2\pi J\left[\omega_{j}(t)\right]\overline{N}\left[\omega_{j}(t)\right]$, where $J\left(\omega\right)$ is the spectral density function and $\overline{N}\left(\omega\right)=1/\left[\exp(\hbar\omega/k_{\rm B}T)-1\right]$ is the average occupation number given by the Bose-Einstein statistics.